# Temporal Smoothness-Aware Rate-Distortion Optimized 4D Gaussian Splatting

### v1

@@ Appendix B Rate-Distortion Performance for Each Scene
Appendix C Analysis on the Number of Pruned Points
To verify that the proposed method effectively reduces redundancy in both dynamic ($\mathcal{G}_{d}$) and static ($\mathcal{G}_{s}$) components, we examine the reduction ratios across different compression levels, as reported in Table 5. The results indicate that at the highest compression setting (Level 1), the reduction rates for dynamic and static points are similarly significant. However, at the lowest compression level (Level 6), the reduction ratio of dynamic points consistently surpasses that of static points. This suggests that when fidelity is prioritized, our method tends to prune dynamic points more aggressively, likely due to the smaller spatial coverage of dynamic points compared to static points. Nonetheless, given that dynamic points typically correspond to salient foreground elements, overly aggressive pruning of dynamic points may adversely impact overall visual quality. This highlights an avenue for future improvement: dynamically adjusting the pruning strategy—for example, by assigning separate weighting coefficients to static and dynamic points within the pruning loss—to better preserve visually salient dynamic components while maintaining high compression efficiency.
@@ Appendix C Analysis on the Number of Pruned Points
Appendix D Visualization of Static and Dynamic Points
We visualize the static and dynamic components separately across different compression levels to further investigate the impact of our proposed compression method, as shown in Figures 8, 9 and 10. Consistent with the quantitative findings discussed in Appendix C, our method tends to prune dynamic points more aggressively than static points, despite dynamic regions generally holding higher visual saliency. Particularly noticeable in the highly compressed results (Level 1) shown in Figures 8 and 9, rapidly moving objects often exhibit noticeable local blur due to aggressive dynamic point pruning. Nevertheless, it’s worth emphasizing that even at moderate compression (Level 6), our method successfully achieves substantial storage savings (approximately 90% on average) while preserving rendering quality across most scenes without significant visual artifacts. This demonstrates our method’s practical balance between storage efficiency and visual fidelity.
@@ Appendix D Visualization of Static and Dynamic Points
Appendix E Analysis of Compressed Representation Size
This appendix details the storage allocation (in MB and percentage) for individual components of our RD-optimized model across various scenes from the Technicolor Table 7 and N3V Table 6 datasets. Results are shown for Low Rate (Level 1) and High Rate (Level 6) settings. These tables provide a granular view of bit distribution for key components such as sorting indices for the masked SH parameters (sort idx), entropy-coded indexes (Indexes), masked wavelet coefficients $\mathbf{F}_{\text{masked}}$, dynamic position displacements ($\boldsymbol{\mu}_{disp}$), static position ($\boldsymbol{\mu}_{0}$), codebooks, logits, and opacity variance parameters ($\beta_{*}^{o}$).
@@ Appendix E Analysis of Compressed Representation Size
Low Rate [$\mathbb{L}$1] High Rate [$\mathbb{L}$6] Low Rate [$\mathbb{L}$1] High Rate [$\mathbb{L}$6]
@@
Low Rate [$\mathbb{L}$1] High Rate [$\mathbb{L}$6] Low Rate [$\mathbb{L}$1] High Rate [$\mathbb{L}$6]
@@
Low Rate [$\mathbb{L}$1] High Rate [$\mathbb{L}$6] Low Rate [$\mathbb{L}$1] High Rate [$\mathbb{L}$6]
@@
Low Rate [$\mathbb{L}$1] High Rate [$\mathbb{L}$6] Low Rate [$\mathbb{L}$1] High Rate [$\mathbb{L}$6]
@@
Low Rate [$\mathbb{L}$1] High Rate [$\mathbb{L}$6] Low Rate [$\mathbb{L}$1] High Rate [$\mathbb{L}$6]
@@
Low Rate [$\mathbb{L}$1] High Rate [$\mathbb{L}$6]

### v2

@@ Appendix C Rate-Distortion Performance for Each Scene
Appendix D Analysis on the Number of Pruned Points
To verify that the proposed method effectively reduces redundancy in both dynamic ($\mathcal{G}_{d}$) and static ($\mathcal{G}_{s}$) components, we examine the reduction ratios across different compression levels, as reported in Table˜7. The results indicate that at the highest compression setting (Level 1), the reduction rates for dynamic and static points are similarly significant. However, at the lowest compression level (Level 6), the reduction ratio of dynamic points consistently surpasses that of static points. This suggests that when fidelity is prioritized, our method tends to prune dynamic points more aggressively, likely due to the smaller spatial coverage of dynamic points compared to static points. Nonetheless, given that dynamic points typically correspond to salient foreground elements, overly aggressive pruning of dynamic points may adversely impact overall visual quality. This highlights an avenue for future improvement: dynamically adjusting the pruning strategy—for example, by assigning separate weighting coefficients to static and dynamic points within the pruning loss—to better preserve visually salient dynamic components while maintaining high compression efficiency.
@@ Appendix D Analysis on the Number of Pruned Points
Appendix E Comparison with Concurrent Work: Light4GS
This section provides a direct comparison with the concurrent work Light4GS liu2025light4gs . To ensure a fair comparison, we retrained our model to match the experimental protocol of Light4GS, which uses a different resolution ($1024\times 768$) and a slightly different subset of the N3V dataset. The results are summarized in Table˜8.
The comparison highlights the different strengths of the two methods. While Light4GS achieves a higher absolute PSNR, our method demonstrates superior performance in terms of compression ratio and rendering speed, particularly in the high-compression setting. Specifically, our high-compression model is more than 2.5x smaller (1.48 MB vs. 3.77 MB) and runs approximately 4x faster ($\sim$160 FPS vs. $\sim$40 FPS).
Furthermore, the PSNR gap is not a fundamental limitation of our compression framework. As demonstrated in our ablation study on higher-fidelity trade-offs (Table˜3), relaxing the rate-loss weight ($\lambda_{R}$) can significantly boost our model’s PSNR. This suggests the performance difference also stems from different design priorities and baseline models (Ex4DGS for ours, 4DGS for Light4GS), rather than the compression methodology itself. Our framework’s primary advantage lies in its flexibility, offering a wide, user-controllable spectrum of rate-distortion-speed trade-offs suitable for diverse deployment scenarios.
Table 8: Direct comparison with Light4GS on the N3V dataset, following their experimental protocol. Our method excels in high-compression and rendering speed, while Light4GS achieves higher peak PSNR.
Model PSNR (dB) Size (MB) FPS
Ours (High-Compression) 27.90 1.48 $\sim$160
Ours (High-Quality) 30.80 12.56 $\sim$100
Light4GS (High-Compression) 31.48 3.77 $\sim$40
Light4GS (High-Quality) 31.69 5.46 $\sim$37
Appendix F Note on Pre-trained Models for Ablation Studies
Our RD-optimized framework follows a two-stage training process: initial pre-training using the Ex4DGS methodology, followed by RD-optimized fine-tuning. For consistency, all primary experimental results reported in this paper (except for Tables˜2 and 3) originate from a single, fixed pre-trained model.
The ablation studies presented in Tables˜2 and 3 were conducted as supplementary experiments. As the original pre-trained model was unavailable at the time of these additional experiments, they utilized a separately pre-trained model. Our methodology involves initiating the RD-optimized fine-tuning after the densification stage of the Ex4DGS schedule, which is prior to its full convergence. Consequently, minor performance variations in the starting point for fine-tuning can occur between different pre-training runs. This explains the slight differences in the baseline performance for the Level 1 and Level 6 configurations between these tables and others in the paper. However, this does not affect our main conclusion: once a pre-trained model is fixed, our framework consistently demonstrates a flexible and effective rate-distortion trade-off controlled by the fine-tuning parameters.
Appendix G Visualization of Static and Dynamic Points
We visualize the static and dynamic components separately across different compression levels to further investigate the impact of our proposed compression method, as shown in Figures˜8, 9 and 10. Consistent with the quantitative findings discussed in Appendix˜D, our method tends to prune dynamic points more aggressively than static points, despite dynamic regions generally holding higher visual saliency. Particularly noticeable in the highly compressed results (Level 1) shown in Figures˜8 and 9, rapidly moving objects often exhibit noticeable local blur due to aggressive dynamic point pruning. Nevertheless, it’s worth emphasizing that even at moderate compression (Level 6), our method successfully achieves substantial storage savings (approximately 90% on average) while preserving rendering quality across most scenes without significant visual artifacts. This demonstrates our method’s practical balance between storage efficiency and visual fidelity.
@@ Appendix G Visualization of Static and Dynamic Points
Appendix H Analysis of Compressed Representation Size
This appendix details the storage allocation (in MB and percentage) for individual components of our RD-optimized model across various scenes from the Technicolor Table˜10 and N3V Table˜9 datasets. Results are shown for Low Rate (Level 1) and High Rate (Level 6) settings. These tables provide a granular view of bit distribution for key components such as sorting indices for the masked SH parameters (sort idx), entropy-coded indexes (Indexes), masked wavelet coefficients $\mathbf{F}_{\text{masked}}$, dynamic position displacements ($\boldsymbol{\mu}_{\text{disp}}$), static position ($\boldsymbol{\mu}_{0}$), codebooks, logits, and opacity variance parameters ($\beta_{*}^{o}$).
@@ Appendix H Analysis of Compressed Representation Size
Low Rate $\mathbb{L}$1] High Rate $\mathbb{L}$6] Low Rate $\mathbb{L}$1] High Rate $\mathbb{L}$6]
@@
Low Rate $\mathbb{L}$1] High Rate $\mathbb{L}$6] Low Rate $\mathbb{L}$1] High Rate $\mathbb{L}$6]
@@
Low Rate $\mathbb{L}$1] High Rate $\mathbb{L}$6] Low Rate $\mathbb{L}$1] High Rate $\mathbb{L}$6]
@@
Low Rate $\mathbb{L}$1] High Rate $\mathbb{L}$6] Low Rate $\mathbb{L}$1] High Rate $\mathbb{L}$6]
@@
Low Rate $\mathbb{L}$1] High Rate $\mathbb{L}$6] Low Rate $\mathbb{L}$1] High Rate $\mathbb{L}$6]
@@
Low Rate $\mathbb{L}$1] High Rate $\mathbb{L}$6]